An oxide of Nitrogen contains 63.1% Oxygen and has a molar mass of 76.0 g/mol


1
Expert's answer
2021-11-15T10:52:04-0500

An oxide of nitrogen contains 63.1% oxygen and has a molar mass of 76.0 g/mol. What is the molecular formula for this compound



Solution


Taking "O=15.9994\quad N=14.0067"


Mass of Oxygen in "100g" sample "= \frac{63.1}{100}×100=63.1g"


Number of moles of Oxygen "=\frac{63.1}{15.9994}=3.9439Moles"


"\%" of Nitrogen "=(100-63.1)"

"=36.9\%"


Mass of Nitrogen in 100g sample "=\frac{36.1}{100}×100=36.1g"


Number of moles of Nitrogen "=\frac{36.9}{14.0067}=2.6345moles"


Moles ratio "N:O=2.6345: 3.9439"


Dividing by smaller number of moles "N:O=1.4970"


Multiplying by 2 and taking integral values

"N:O=2:3"

Empirical formula is "N_2O_3"

Molar mass of "N_2O_3=(14.0067\u00d72)+(15.9994\u00d73)"


"=76.0116g"

Relative Molecular Mass "=76" g/mole

Molecular formula "=N_2O_3"